# Stable Trajectory Clustering: An Efficient Split and Merge Algorithm

### v1

@@ 6 $W_{temp}\{i\}\leftarrow C_{i}^{R_{i}};$
9 for j from 1 to $W_{\text{move}}-1$ do $\mathcal{B}_{\text{sim}}\leftarrow(W_{\text{temp}}\{j\}=W_{\text{temp}}\{j+1\})\;?\;1:0$;
10 end for
[…]
12 foreach $i\in\{i^{\prime}\mid\mathcal{B}_{\text{sim}}[i^{\prime}]=0\}$ do $W_{s}\leftarrow(i-1)\times S+1;$ $W_{e}\leftarrow W_{s}+W;$
13 $p\leftarrow[W_{s},W_{e}];$
14 $C_{s}.\text{add}(C_{p}^{R_{p}});$ end foreach
[…]
16 for i from 1 to $|\mathcal{I}_{\text{sim}}|$ do $e\leftarrow$lastElement$(\mathcal{I}_{\text{sim}}\{i\})$;
@@ 19 $p\leftarrow[W_{s},W_{e}];$
Algorithm 5 Sub-trajectory clusteirng
The detailed steps of this algorithm are as follows: Firstly, based on the parameters $S$ and $W$, the trajectories within each window are clustered (lines 1-8). Then, each pair of continuous windows should be compared (lines 10 and 11). Continuous ranges with identical results are therefore merged and identified as a single range in which the sub-trajectories all have the same patterns (lines 13-23). Both whole-trajectory and sub-trajectory clustering follow the same procedures, utilizing the results of the split and merge events of line segment clustering in each time interval. Nevertheless, while the whole-trajectory clustering algorithm focuses on patterns and similarities across entire trajectories, sub-trajectory clustering examines specific parts of trajectories via a sliding window.
V Stable Trajectory Clustering Algorithm
@@
Definition V.1 (Stable trajectory clustering).
@@
To further enhance the results of trajectory clustering, we can evaluate the outlier trajectories, which are the trajectories of moving objects that deviate from their respective clusters for a few intervals and create anomalies. Such anomalies may be caused by noise in GPS trackers [8] or conditions that are dependent on the nature of the object’s movement. Sometimes minor deviations are often inconsequential and can be disregarded. Whether or not deviations should be ignored depends on the number of time intervals over which the objects’ trajectories remain split from their clusters. Additionally, the magnitude of their deviation from the clusters within these time intervals is important. Indeed, we aim to determine whether, despite the anomaly causing a trajectory to split from its cluster temporarily, it still belongs to the same cluster, or if this anomaly represents a genuine split. As an example, the trajectories of six moving objects from top to bottom in different colors are shown in Fig. 2(a). There are two vertical and horizontal crossed corridors in this example. In the middle, due to the anomalies, the sub-trajectory clustering and whole-trajectory clustering have detected five clusters and three clusters, respectively (Fig. 2(b) and Fig. 2(c)). Now, if the objects spend significantly more time traversing the vertical corridor compared to the horizontal one, the clustering results may be different. In fact, the stability of the trajectory can smooth the impact of movements in the horizontal corridor, and different results may be achieved as shown in Fig. 2(d) and Fig. 2(e), respectively.
@@
Figure 2: Trajectory clustering based on the concept of stability. (a) The trajectories of six moving objects from top to bottom. (b) Detection of five clusters by sub-trajectory clustering. (c) Identification of three clusters by whole-trajectory clustering. (d) Disregarding the mid-trajectory anomalies and considering three sub-trajectories within a single cluster. (e) Ignoring the mid-trajectory anomalies completely and considering all trajectories as belonging to one cluster.
@@
Building upon the discussions of whole-trajectory clustering and sub-trajectory clustering presented earlier, we can now obtain trajectory clusters and outliers. Then, the outlier or split trajectories can be assessed to determine whether their split should be considered a real split or not. Suppose $C_{num}$ represents a cluster of trajectories and $\pi_{a}$, $\pi_{b}$ are two outlier or split trajectories which, due to the anomalies in some time intervals, do not belong to the cluster $C_{num}$. Thus, they should be evaluated by the stable trajectory clustering algorithm. In the following, we will begin this evaluation with $\pi_{a}$. We intend to determine the distance between the modified outlier trajectory of $\pi_{a}$ that is denoted as $\pi^{\prime}_{a}$ and the members of cluster $C_{num}$ such as $\pi_{c}$ using definition III.4. To this aim, suppose $l_{\pi_{a}}$ represents a line segment of $\pi_{a}$ and $l_{\pi_{c}}$ is a line segment of $\pi_{c}$ in $lc^{R_{i}}$ which is the current line segment cluster in $i$-th time interval. For these two outlier trajectories, two states may occur for their line segments in this time interval:
@@
Figure 3: Two states for outlier trajectories $\pi_{a}$ and $\pi_{b}$. The trajectories in black are members of cluster $C_{num}$. The dashed line segments represent the approximate movement of objects that have deviated from their respective line segment clusters.
Suppose the first state mentioned occurs within this time interval; in that case, it is sufficient to calculate the average Euclidean distance between $l_{\pi_{a}}$ and $l_{\pi_{c}}$ directly. This is because both line segments are in the same line segment cluster and are neighbors (definitions III.7 to III.9); otherwise, the second state occurs and the amount by which the distance between $l_{\pi_{a}}$ and $lc^{R_{i}}$ needs to be reduced for $l_{\pi_{a}}$ to become a member of $lc^{R_{i}}$ should be determined. Consider $l_{N}$ as a borderline segment (definition III.10) in $lc^{R_{i}}$. It is evident that the average Euclidean distance between $l_{\pi_{a}}$ and its nearest line segment $l_{N}$ exceeds $\varepsilon$ by $\delta_{i}$. This is because the distance being less than or equal to $\varepsilon$ implies that $l_{\pi_{a}}$ and $l_{N}$ are in the same cluster and are neighbors. Here, $\delta_{i}$ is the deviation in the $i$-th time interval. Accordingly, the average Euclidean distance between $l_{\pi_{a}}$ and $l_{\pi_{c}}$ in the current line segment cluster of the $i$-th time interval should be reduced by $\delta_{i}$. Thus, line segment $l_{\pi_{a}}$ is modified as $l^{\prime}_{\pi_{a}}$. This modified line segment shows the most likely approximate movement of object $a$ in this time interval. Therefore, the piecewise function for obtaining the set of average Euclidean distances in $T-1$ time intervals is defined as follows:
$D(\pi^{\prime}_{a},\pi_{c})_{i}^{R_{i}}=\begin{cases}dist(l_{\pi_{a}},l_{\pi_{c}}),&lc^{R_{i}}(l_{\pi_{a}})=lc^{R_{i}}(l_{\pi_{c}}),\\ dist(l_{\pi_{a}},l_{\pi_{c}})-\delta_{i},&o.w,\\ \end{cases}$ (5)
where $i$ takes values in the range from $1$ to $T-1$ and $\delta_{i}=l_{N}-\varepsilon$. This process is shown in Fig. 4. The maximum value of these distances represents their overall average Euclidean distance and is denoted as $M_{(\pi^{\prime}_{a},\pi_{c})}$. It indicates the maximum possible distance at which the outlier trajectory $\pi_{a}$ would be considered a member of cluster $C_{num}$, disregarding its anomalies.
$M_{(\pi^{\prime}_{a},\pi_{c})}=\max\{D(\pi^{\prime}_{a},\pi_{c})_{i}^{R_{i}}\quad|\quad i=1,\ldots,T-1\}.$ (6)
@@
$\mu_{min}=\min\left\{\displaystyle M_{(\pi^{\prime}_{i},\pi_{j})}\;\middle|\;i=1,\ldots,N_{O},\;j=1,\ldots,N_{c}\right\}.$ (7)
Figure 4: The average Euclidean distance between $l_{\pi_{a}}$ and $l_{\pi_{c}}$ with the assistance of the nearest line segment $l_{N}$ in current line segment cluster. $l_{\pi_{a}^{\prime}}$ is denoted as the approximate movement of the object in the $i$-th time interval.
As mentioned in definition V.1, the stable trajectory clustering algorithm aims to assess whether it is possible to ignore the anomalies that caused an outlier trajectory split from its respective cluster or not. The number of time intervals and the deviation amount in these intervals are important factors for deciding whether or not to ignore anomalies. Specifically, one of the following four cases may occur:
Case 1.
@@
Definition V.2 (Mean absolute deviation).
@@
$RMD=\sum_{RD>\mu_{min}}\absolutevalue{RD-\mu_{min}},$ (9)
$LMD=\sum_{LD<\mu_{min}}\absolutevalue{LD-\mu_{min}}.$ | .
Taking into account the aforementioned formula, and based on the fact that the average Euclidean distance of an outlier trajectory to cluster $C_{num}$ must be $\mu_{min}$ for it to be considered a member, if $RMD$ is greater than $LMD$, then a real split has occurred, and the deviations cannot be ignored. Specifically, this occurs when the deviation in the time intervals has been very significant (Case 1. or Case 2.) or the number of time intervals in which the trajectory deviates from its cluster is more than the number of time intervals in which the trajectory is in its cluster (Case 3.); in contrast, if RMD is less than LMD, then, despite the anomalies, they belong to the same cluster and are stable in cluster $C_{num}$ (Case 4.). The stable trajectory clustering is presented in Algorithm 6. This method improves trajectory clustering results by evaluating outlier trajectories and their association with existing clusters. It begins by checking if there are any outlier trajectories. If outliers exist, the algorithm iterates through each cluster, comparing each trajectory’s line segments with those of other trajectories, and calculating distances and deviations. It identifies the maximum distance among all these deviations and sets a minimum threshold, $\mu_{min}$. The algorithm further refines the cluster assignments by computing real distances and evaluating right and left deviations, deciding which outlier trajectories can remain in the respective cluster despite their deviations. If any cluster assignments are altered, it updates the clusters and returns the new set; otherwise, it returns the original set.
@@
1 if $\mathcal{O}=\emptyset$ then return $C$ end if
2foreach cluster $c_{k}\in C$ do $\mathcal{M}_{k}\leftarrow c_{k}.$allMembers;
3 $\mathcal{O}_{k}\leftarrow$ assignOutliers$(c_{k},\mathcal{O})$;
@@ 4 $\mu_{\min}\leftarrow\infty$;
5 foreach $\pi_{m}\in\mathcal{M}_{k},\pi_{o}\in\mathcal{O}_{k}$ do $M_{(\pi_{m},\pi_{o})}\leftarrow-\infty$;
6 for $t$ from 1 to $T-1$ do
7if $lc^{R_{t}}(l_{\pi_{m}})=lc^{{R_{t}}}(l_{\pi_{o}})$ then $D(\pi_{m},\pi_{o})_{t}^{R_{t}}\leftarrow\text{dist}(l_{\pi_{m}},l_{\pi_{o}})$; else $\delta\leftarrow\text{nearestDistance}(\pi_{o}^{t},\mathcal{M}_{k}^{t})-\varepsilon$;
8 $D(\pi_{m},\pi_{o})_{t}^{R_{t}}\leftarrow\text{dist}(l_{\pi_{m}},l_{\pi_{o}})-\delta$; end if
[…]
11end foreach
12foreach $\pi_{o}\in\mathcal{O}_{k}$ do foreach $\pi_{m}\in\mathcal{M}_{k}$ do $\text{LMD}\leftarrow 0$, $\text{RMD}\leftarrow 0$;
13 for $t$ from 1 to $T-1$ do $D(\pi_{m},\pi_{o})_{t}^{R_{t}}\leftarrow\text{dist}(l_{\pi_{m}},l_{\pi_{o}})$;
14 if $D(\pi_{m},\pi_{o})_{t}^{R_{t}}<\mu_{\min}$ then $\text{LMD}\leftarrow\text{LMD}+(\mu_{\min}-D(\pi_{m},\pi_{o})_{t}^{R_{t}})$; else $\text{RMD}\leftarrow\text{RMD}+(D(\pi_{m},\pi_{o})_{t}^{R_{t}}-\mu_{\min})$; end if
15end for
[…]
17 break; end if
18end foreach
19
20 end foreach
@@ 8 $D(\pi_{m},\pi_{o})_{t}^{R_{t}}\leftarrow\text{dist}(l_{\pi_{m}},l_{\pi_{o}})-\delta$; end if
The details of this algorithm are as follows: Initially, a function is defined to identify clusters and respective outlier trajectories (lines 3 and 4). Afterwards, an iteration is performed for each cluster and its outlier trajectories. This iteration tries to use the mean absolute deviation concept to find the outlier trajectories that, despite their anomalies, still belong to their respective clusters (lines 7- 28). Finally, if the clusters change, they will be updated and returned as a new set (line 29). The stable trajectory clustering algorithm can be utilized to analyze outlier trajectories, which due to their anomalies, are split from their respective cluster in some time intervals. In this analysis, it is important to consider both the number of time intervals in which deviations occur and the magnitudes of those deviations. We therefore used the concept of mean absolute deviation to decide if, despite the anomalies, outlier trajectories could be assigned to their nearest cluster. The stable trajectory clustering algorithm can be applied to both whole-trajectory clustering and sub-trajectory clustering problems.
VI Empirical Results

### v2

@@ 6 $p\leftarrow[W_{s},W_{e}];$
8 $W_{temp}\{i\}\leftarrow C_{i}^{R_{i}};$
9
[…]
11 for j from 1 to $W_{\text{move}}-1$ do
12 $\mathcal{B}_{\text{sim}}\leftarrow(W_{\text{temp}}\{j\}=W_{\text{temp}}\{j+1\})\;?\;1:0$;
13
[…]
15 foreach $i\in\{i^{\prime}\mid\mathcal{B}_{\text{sim}}[i^{\prime}]=0\}$ do
16 $W_{s}\leftarrow(i-1)\times S+1;$ $W_{e}\leftarrow W_{s}+W;$
@@ 17 $p\leftarrow[W_{s},W_{e}];$
20 for i from 1 to $|\mathcal{I}_{\text{sim}}|$ do
21 $e\leftarrowlastElement$$(\mathcal{I}_{\text{sim}}\{i\})$;
@@ 24 $p\leftarrow[W_{s},W_{e}];$
Algorithm 5 Sub-trajectory clustering
The detailed steps of this algorithm are as follows: Firstly, based on the parameters $S$ and $W$, the trajectories within each window are clustered (lines 1-8). Then, each pair of continuous windows should be compared (lines 10 and 11). Continuous ranges with identical results are therefore merged and identified as a single range in which the sub-trajectories all have the same patterns (lines 13-23). The time complexity is $\mathcal{O}\left(\frac{m}{S}\cdot(n^{2}\cdot W)\right)$, because the algorithm repeats the whole-trajectory clustering process on a window of size $W$ for approximately $m/S$ steps.
Both whole-trajectory and sub-trajectory clustering follow the same procedures, utilizing the results of the split and merge events of line segment clustering in each time interval. Nevertheless, while the whole-trajectory clustering algorithm focuses on patterns and similarities across entire trajectories, sub-trajectory clustering examines specific parts of trajectories via a sliding window.
5. Stable Trajectory Clustering Algorithm
@@
Definition 5.0 (Stable trajectory clustering).
@@
To further enhance the results of trajectory clustering, we can evaluate the outlier trajectories, which are the trajectories of moving objects that deviate from their respective clusters for a few intervals and create anomalies. Such anomalies may be caused by noise in GPS trackers (20) or conditions that are dependent on the nature of the object’s movement. Sometimes minor deviations are often inconsequential and can be disregarded. Whether or not deviations should be ignored depends on the number of time intervals over which the objects’ trajectories remain split from their clusters. Additionally, the magnitude of their deviation from the clusters within these time intervals is important. Indeed, we aim to determine whether, despite the anomaly causing a trajectory to split from its cluster temporarily, it still belongs to the same cluster, or if this anomaly represents a genuine split.
As an example, the trajectories of six moving objects from top to bottom in different colors are shown in Fig. 2(a). There are two vertical and horizontal crossed corridors in this example. Due to the mid-trajectory deviations where objects briefly enter the horizontal corridor, the sub-trajectory clustering and whole-trajectory clustering produce fragmented results, and split the trajectories into five and three clusters, respectively (Fig. 2(b) and Fig. 2(c)). In contrast, the stability of the trajectory can mitigate the impact of movements in the horizontal corridor, and different results may be achieved, as shown in Figs. 2(d) and 2(e), respectively.
@@
Figure 2. Trajectory clustering based on the concept of stability. (a) The trajectories of six moving objects from top to bottom. (b) Detection of five clusters by sub-trajectory clustering. (c) Identification of three clusters by whole-trajectory clustering. (d) Disregarding the mid-trajectory anomalies and considering three sub-trajectories within a single cluster. (e) Ignoring the mid-trajectory anomalies completely and considering all trajectories as belonging to one cluster.
@@
Building upon the discussions of whole-trajectory clustering and sub-trajectory clustering presented earlier, we can now obtain trajectory clusters and outliers. Then, the outlier or split trajectories can be assessed to determine whether their split should be considered a real split or not. Suppose $C_{num}$ represents a cluster of trajectories and $\pi_{a}$, $\pi_{b}$ are two outlier or split trajectories which, due to the anomalies in some time intervals, do not belong to the cluster $C_{num}$. Thus, they should be evaluated by the stable trajectory clustering algorithm. In the following, we will begin this evaluation with $\pi_{a}$. We intend to determine the distance between the modified outlier trajectory of $\pi_{a}$ that is denoted as $\pi^{\prime}_{a}$ and the members of cluster $C_{num}$ such as $\pi_{c}$ using definition 3.4. To this aim, suppose $l_{\pi_{a}}$ represents a line segment of $\pi_{a}$ and $l_{\pi_{c}}$ is a line segment of $\pi_{c}$ in $lc^{R_{i}}$ which is the current line segment cluster in $i$-th time interval. For these two outlier trajectories, two states may occur for their line segments in this time interval:
@@
Figure 3. Two states for outlier trajectories $\pi_{a}$ and $\pi_{b}$. The trajectories in black are members of cluster $C_{num}$. The dashed line segments represent the approximate movement of objects that have deviated from their respective line segment clusters.
Suppose the first state mentioned occurs within this time interval; in that case, it is sufficient to calculate the average Euclidean distance between $l_{\pi_{a}}$ and $l_{\pi_{c}}$ directly. This is because both line segments are in the same line segment cluster and are neighbors (definitions 3.8 to 3.10); otherwise, the second state occurs and the amount by which the distance between $l_{\pi_{a}}$ and $lc^{R_{i}}$ needs to be reduced for $l_{\pi_{a}}$ to become a member of $lc^{R_{i}}$ should be determined. Consider $l_{N}$ as a borderline segment (definition 3.11) in $lc^{R_{i}}$. It is evident that the average Euclidean distance between $l_{\pi_{a}}$ and its nearest line segment $l_{N}$ exceeds $\varepsilon$ by $\delta_{i}$. This is because the distance being less than or equal to $\varepsilon$ implies that $l_{\pi_{a}}$ and $l_{N}$ are in the same cluster and are neighbors. Here, $\delta_{i}$ is the deviation in the $i$-th time interval. Accordingly, the average Euclidean distance between $l_{\pi_{a}}$ and $l_{\pi_{c}}$ in the current line segment cluster of the $i$-th time interval should be reduced by $\delta_{i}$. Thus, line segment $l_{\pi_{a}}$ is modified as $l^{\prime}_{\pi_{a}}$. This modified line segment shows the most likely approximate movement of object $a$ in this time interval. Therefore, the piecewise function for obtaining the set of average Euclidean distances in T1T-1 time intervals is defined as follows:
[…]
where $i$ takes values in the range from $1$ to T1T-1 and $\delta_{i}=l_{N}-\varepsilon$. This process is shown in Fig. 4. The maximum value of these distances represents their overall average Euclidean distance and is denoted as $M_{(\pi^{\prime}_{a},\pi_{c})}$. It indicates the maximum possible distance at which the outlier trajectory $\pi_{a}$ would be considered a member of cluster $C_{num}$, disregarding its anomalies.
(6) $M_{(\pi^{\prime}_{a},\pi_{c})}=\max\{D(\pi^{\prime}_{a},\pi_{c})_{i}^{R_{i}}\quad|\quad i=1,\ldots,T-1\}.$
@@
(7) $\mu_{min}=\min\left\{\displaystyle M_{(\pi^{\prime}_{i},\pi_{j})}\;\middle|\;i=1,\ldots,N_{O},\;j=1,\ldots,N_{c}\right\}.$
Figure 4. The average Euclidean distance between $l_{\pi_{a}}$ and $l_{\pi_{c}}$ with the assistance of the nearest line segment $l_{N}$ in current line segment cluster. $l_{\pi_{a}^{\prime}}$ is denoted as the approximate movement of the object in the $i$-th time interval.
As mentioned in definition 5.1, the stable trajectory clustering algorithm aims to assess whether it is possible to ignore the anomalies that caused an outlier trajectory split from its respective cluster or not. The number of time intervals and the deviation amount in these intervals are important factors for deciding whether or not to ignore anomalies. Specifically, one of the following four cases may occur:
Case 1.:
@@
Definition 5.0 (Mean absolute deviation).
@@
(9) $|,\displaystyle RMD=\sum_{RD>\mu_{min}}\absolutevalue{RD-\mu_{min}}$,
$|.\displaystyle LMD=\sum_{LD<\mu_{min}}\absolutevalue{LD-\mu_{min}}$.
Taking into account the aforementioned formula, and based on the fact that the average Euclidean distance of an outlier trajectory to cluster $C_{num}$ must be $\mu_{min}$ for it to be considered a member, if $RMD$ is greater than $LMD$, then a real split has occurred, and the deviations cannot be ignored. Specifically, this occurs when the deviation in the time intervals has been very significant (Case 1. or Case 2.) or the number of time intervals in which the trajectory deviates from its cluster is more than the number of time intervals in which the trajectory is in its cluster (Case 3.); in contrast, if RMD is less than LMD, then, despite the anomalies, they belong to the same cluster and are stable in cluster $C_{num}$ (Case 4.).
The stable trajectory clustering is presented in Algorithm 6. This method improves trajectory clustering results by evaluating outlier trajectories and their association with existing clusters. It begins by checking if there are any outlier trajectories. If outliers exist, the algorithm iterates through each cluster, comparing each trajectory’s line segments with those of other trajectories, and calculating distances and deviations. It identifies the maximum distance among all these deviations and sets a minimum threshold, $\mu_{min}$. The algorithm further refines the cluster assignments by computing real distances and evaluating right and left deviations, deciding which outlier trajectories can remain in the respective cluster despite their deviations. If any cluster assignments are altered, it updates the clusters and returns the new set; otherwise, it returns the original set.
@@
1 if $\mathcal{O}=\emptyset$ then
2 return $C$
3foreach cluster $c_{k}\in C$ do
4 $\mathcal{M}_{k}\leftarrow c_{k}.$allMembers;
[…]
@@ 6 $\mu_{\min}\leftarrow\infty$;
7 foreach $\pi_{m}\in\mathcal{M}_{k},\pi_{o}\in\mathcal{O}_{k}$ do
8    $M_{(\pi_{m},\pi_{o})}\leftarrow$-$\infty$;
9    for $t$ from 1 to T1T-1 do
10
11       if $lc^{R_{t}}(l_{\pi_{m}})=lc^{{R_{t}}}(l_{\pi_{o}})$ then
12          $D(\pi_{m},\pi_{o})_{t}^{R_{t}}\leftarrow\text{dist}(l_{\pi_{m}},l_{\pi_{o}})$;
13       else
14          $\delta\leftarrow\text{nearestDistance}(\pi_{o}^{t},\mathcal{M}_{k}^{t})-\varepsilon$;
15          $D(\pi_{m},\pi_{o})_{t}^{R_{t}}\leftarrow\text{dist}(l_{\pi_{m}},l_{\pi_{o}})-\delta$;
16       $M_{(\pi_{m},\pi_{o})}\leftarrow\max(M_{(\pi_{m},\pi_{o})},D(\pi_{m},\pi_{o})_{t}^{R_{t}})$;
17    $\mu_{\min}\leftarrow\min(\mu_{\min},M_{(\pi_{m},\pi_{o})})$;
18foreach $\pi_{o}\in\mathcal{O}_{k}$ do
19 foreach $\pi_{m}\in\mathcal{M}_{k}$ do
20    $\text{LMD}\leftarrow 0$, $\text{RMD}\leftarrow 0$;
21    for $t$ from 1 to T1T-1 do
22       $D(\pi_{m},\pi_{o})_{t}^{R_{t}}\leftarrow\text{dist}(l_{\pi_{m}},l_{\pi_{o}})$;
23       if $D(\pi_{m},\pi_{o})_{t}^{R_{t}}<\mu_{\min}$ then
24          $\text{LMD}\leftarrow\text{LMD}+(\mu_{\min}-D(\pi_{m},\pi_{o})_{t}^{R_{t}})$;
25       else
26          $\text{RMD}\leftarrow\text{RMD}+(D(\pi_{m},\pi_{o})_{t}^{R_{t}}-\mu_{\min})$;
27    if $\text{LMD}>\text{RMD}$ then
28       $c_{k}\leftarrow c_{k}\cup\{\pi_{o}\}$;
29         break;
@@ 28       $c_{k}\leftarrow c_{k}\cup\{\pi_{o}\}$;
The details of this algorithm are as follows: Initially, a function is defined to identify clusters and respective outlier trajectories (lines 3 and 4). Afterwards, an iteration is performed for each cluster and its outlier trajectories. This iteration tries to use the mean absolute deviation concept to find the outlier trajectories that, despite their anomalies, still belong to their respective clusters (lines 7- 28). Finally, if the clusters change, they will be updated and returned as a new set (line 29). The computational cost is $\mathcal{O}(|O_{k}|\cdot|c_{k}|\cdot m)$, as the algorithm evaluates the distance between every outlier and the cluster core members across all time intervals.
The stable trajectory clustering algorithm can be utilized to analyze outlier trajectories, which due to their anomalies, are split from their respective cluster in some time intervals. In this analysis, it is important to consider both the number of time intervals in which deviations occur and the magnitudes of those deviations. We therefore used the concept of mean absolute deviation to decide if, despite the anomalies, outlier trajectories could be assigned to their nearest cluster. The stable trajectory clustering algorithm can be applied to both whole-trajectory clustering and sub-trajectory clustering problems.
6. Empirical Results